OpenStudy (pcywife):

if a b and c are positive integers such that square root (a/b) = 8 c and ac = b, what is the value of c

OpenStudy (welshfella):

if we plug b = ac into sqrtr(a/b) we get sqrt (a /ac0 = sqrt (1 / c) so we can say sqrt (1/c) = 8c solve for c

OpenStudy (welshfella):

square both sides 1 / c = 64c^2

OpenStudy (welshfella):

that will not give a positive integer though. Are you sure you have the question right?